solve the following system by any method: x=7/4y+8 4x-7y=32
Answer by Alan3354(69443)

solve the following system by any method: x=7/4y+8 4x-7y=32
------------------
x=7/4y+8 eqn 1 (assuming you mean x = (7/4)y + 8
4x-7y=32
Multiply eqn 1 by 4
4x = 7y+32
4x-7y = 32 This is eqn 2 again, so it's one eqn is 2 unknowns.
This is a straight line. Any point on it satisfies this one equation.